OpenStudy (anonymous):

evaluate the expression at sqrt(-5)sqrt(-20) writing the result in a+bi form

Directrix (directrix):

Is this √(-5) * √-20) ?

OpenStudy (anonymous):

\[\sqrt{-5}\sqrt{-20}\]

OpenStudy (anonymous):

ya thats what i thought too thats why i asked cause it says its wrong so idk

Directrix (directrix):

i√5 times 2i√5 = 2i² times 5 = 2(-1) 5 = -10 = - 10 + 0i

OpenStudy (anonymous):

ha ha i guess i did to casue i tried entering that answer but thanks

Directrix (directrix):

Are you saying that -10 + 0i is wrong?

OpenStudy (anonymous):

oh no i was talking about the first answer you put

Directrix (directrix):

Okay.